Team A has 7 male and 3 female players, and team B has 6 male and 4 female players.Two players are to be selected from each of the two teams. in how many ways can the four players be selected so that at least one of them is female?
Answer by sudhanshu_kmr(1152)


Total number of ways to select 4 players = 10C2 * 10C2 = 45*45 = 2025

numbers of ways when all are male = 7C2 * 6C2 = 21*15 = 315

number of ways when at least one of them is female = 2025 - 315 = 1710

if any doubt, you are welcome to contact.